Which of the following is the indefinite integral of x2 + 7?




Select one:
A. x3 + 7x + c
B. 2x + c
C. (1/3)x3 + 7x + c
D. ½ x3 + 7x + c
1
Expert's answer
2020-10-27T18:46:20-0400

"\intop(x^2+7)dx=\intop(x^2)dx+7\intop(dx)=x^3/3+7x+const"

the integral was presented as the sum of the table integrals. The option C is correct.